OpenStudy (brooke..help00):

@lowkey.s @asc.bchs Below are the steps to solve an equation: Step 1: |x − 2| + 3 = 7 Step 2: |x − 2| = 7 − 3 Step 3: |x − 2| = 4 Which of the following is a correct next step to solve the equation? x + 2 = −4 −x − 2 = 4 x + 2 = 4 x − 2 = −4

OpenStudy (lowkey.s):

Any guess?

OpenStudy (brooke..help00):

the first one

OpenStudy (lowkey.s):

Take another try

OpenStudy (lowkey.s):

@brooke..help00

OpenStudy (brooke..help00):

Oh the third one I didn't realize that said negative 4

OpenStudy (lowkey.s):

no

OpenStudy (lowkey.s):

It's alright but, lets figure this out

OpenStudy (brooke..help00):

ok:)

OpenStudy (brooke..help00):

So what do I do?

OpenStudy (mathmale):

Solve the problem yourself in its entirety. If you can do that, you should be able to identify the correct next step.

OpenStudy (brooke..help00):

It has to be the last one

OpenStudy (brooke..help00):

Hello? Anyone?

zepdrix (zepdrix):

\[\large\rm |x-2|=4\] Hey Brooke :) We can drop the absolute bars by putting a plus/minus symbol on our expression.\\[\large\rm \pm(x-2)=4\]Sometimes we write the plus/minus on the other side because it's more convenient.\[\large\rm x-2=\pm4\]This corresponds to two equations,\[\large\rm x-2=4\]\[\large\rm x-2=-4\]

OpenStudy (brooke..help00):

Thank you so much!!!!!!!!!!!!!!!!

OpenStudy (lowkey.s):

Oh wait its B... its asking the correct next step Neverming it is B

zepdrix (zepdrix):

If we had left the plus/minus on the left side of the equation, then,\[\large\rm -(x-2)=4\]could have been a correct next step. But notice that is equivalent to \(\large\rm -x+2=4\) after districuting the negative, which is not option B, ya?

zepdrix (zepdrix):

distributing*

zepdrix (zepdrix):

Ya it was D :P

OpenStudy (lowkey.s):

Oh

OpenStudy (lowkey.s):

Okay cause I thought it was D but, than I had a second thought so, zepdrix is right